According to a study, 69% o all males between the ages of 18 and 24 live at home. Unmarried college st de sliving in a mare count a ving athome. Suppose that a survey is administered and 166 of 234 respondents indicated that they live at home. (a) Use the normal approximation to the binomial to approximate the probability that at least 166 respondents live at home? (b) Do the results from part (a) contradict the study? (a) P(X2 166)(Round to four decimal places as needed.)
Answer #1

Sample Proportion = 166/234

= 0.7094

P(X>=166)

= P(Z>(\frac{0.7094-0.69}{\sqrt{\frac{0.69*0.31}{234}}}))

P(Z>0.64)

Use excel function

1-NORMSDIST(0.64)

=0.2605